Two tiny conducting spheres are identical and carry charges of -15.9uC and +64.3uC. They are separated by a distance of 2.44 cm. (a) What is the magnitude of the force that each sphere experiences? (b) The spheres are brought into contact and then separated to a distance of 2.44 cm. Determine the magnitude of the force that each sphere now experiences.

Thanks for the help!
Answer #1

a) for the magnitude of the force

magnitude of force = k * q1 * q2/d^2

magnitude of force = 9 *10^9 * (15.9 *10^-6 * 64.3 *10^-6)/0.0244^2

magnitude of force = 15455 N

b) final charge after touching

q = (-15.9 + 64.3)/2 = 24.2 uC

magnitude of force = 9 *10^9 * (24.2 *10^-6)^2/0.0244^2

magnitude of force = 8853.0 N